OpenStudy (anonymous):

find arc length of a circle with a radius of 5 and a central angle of 7 rad.

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

 \[\Large \text{Arc length} = \frac{\text{radian angle}}{2\pi \ \text{radians}}*(\text{circumference})\] \[\Large S = \frac{\theta}{2\pi}*(2\pi*r)\] \[\Large S = \frac{\theta}{\cancel{2\pi}}*(\cancel{2\pi}*r)\] \[\Large S = \theta*r\]

jimthompson5910 (jim_thompson5910):

The formula \(\Large S = \theta*r\) works only if \(\Large \theta\) (greek letter theta) is in radians

OpenStudy (anonymous):

so 35? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

hmm something is off though, I think theta needs to be restricted

jimthompson5910 (jim_thompson5910):

the formula only works if theta is between 0 and 2pi